OpenStudy (anonymous):

what is the LCM of 14, 21, 28 i got 28 is ti correct?

OpenStudy (amistre64):

lets see, 21 times what is 28?

OpenStudy (anonymous):

it can't

OpenStudy (amistre64):

14 = 7*2 21 = 7*3 28 = 7*2*2 then lets split these into prime factors and see what we can make 14 = 7*2 21 = 7*3 28 = 7*2*2 14 needs a 2 and a 3; 21 needs a 2 and a 2; and 28 needs a 3 such that LCM = 2*2*3*7 most likely

OpenStudy (amistre64):

the double 2 might be a bit much ... its been a while since i tried thinking of these

OpenStudy (anonymous):

this what i did: i just listed the multiples and found the least one which was 28

OpenStudy (amistre64):

LCM = 2*3*7 = 42 ; and 28 doesnt multiply to get 42 ... looks like we would need it 42*2 = 84, not 28

OpenStudy (amistre64):

14: 28,42,56,70,84 21: 42, 63, 84 28: 56, 84

OpenStudy (amistre64):

the smallest multiple they have in common is 84